OpenStudy (anonymous):

(-3x^4 y^(-5))(4x^(-3) y^2)

OpenStudy (calculusfunctions):

Simplify\[(-3x ^{4}y ^{-5})(4x ^{-3}y ^{2})\]Is this your question?

OpenStudy (anonymous):

yes it is

OpenStudy (anonymous):

how do i solve this problem

OpenStudy (calculusfunctions):

Do you know your exponent laws?

OpenStudy (calculusfunctions):

@jmays14 it's all multiplication. But first do you know the exponent law which states that when you multiply two powers with the same base, you add their exponents?

OpenStudy (anonymous):

no i dont could you help me out

OpenStudy (calculusfunctions):

Sure. The product law states that when you multiply two powers with the same base, you add their exponents. e.g.\[x ^{a}\times x ^{b}=x ^{a +b}\]Understand?

OpenStudy (anonymous):

so im going to multiply 3 and 4

OpenStudy (anonymous):

how would i show my work

OpenStudy (calculusfunctions):

Step 1: (-3)(4) = ?

OpenStudy (anonymous):

-12

OpenStudy (calculusfunctions):

Step 2:\[x ^{4}\times x ^{-3}=?\]

OpenStudy (anonymous):

7

OpenStudy (calculusfunctions):

That's a -3. The feature here makes the exponents 2 and 3 look alike.

OpenStudy (calculusfunctions):

No,! Remember you add the exponents. Try again.

OpenStudy (anonymous):

2

OpenStudy (calculusfunctions):

\[x ^{4}\times x ^{-3}=x ^{4+(-3)}=?\]

OpenStudy (calculusfunctions):

No, I told you that its a -3, not a -2. Please pay attention!!!

OpenStudy (anonymous):

1

OpenStudy (calculusfunctions):

so you mean x, right?

OpenStudy (anonymous):

yea

OpenStudy (calculusfunctions):

Step 3:\[y ^{-5}\times y ^{2}=?\]This time that exponent is 2.

OpenStudy (anonymous):

-7

OpenStudy (calculusfunctions):

\[y ^{-5}\times y ^{2}=y ^{-5+2}=?\]

OpenStudy (anonymous):

-3

OpenStudy (calculusfunctions):

you mean\[y ^{-3}\]right?

OpenStudy (anonymous):

yes now how do i put it all together

OpenStudy (calculusfunctions):

Now when you multiply the results of all three steps, what do you get?

OpenStudy (anonymous):

how do you put it all together

OpenStudy (calculusfunctions):

I just told you, multiply the results of all three steps.

OpenStudy (anonymous):

i got 12 can you right it out for me so ill kno how to do the next 1

OpenStudy (calculusfunctions):

The answer is\[-12xy ^{-3}\]or\[\frac{ -12x }{ y ^{3} }\] Remember that exponent that looks like a 2, is actually a 3.

OpenStudy (calculusfunctions):

Understand?